OpenStudy (coralskates):

Solve for x: 3|x − 3| + 2 = 14 No solutions x = −1, x = 8.3 x = 0, x = 7 x = −1, x = 7

OpenStudy (sushi121212):

The absolute sign means that if you get a negative inside the sign, it's a positive, and if the number is positive inside the sign, it is still positive. 3|x − 3| + 2 = 14 You can plug in the answer choices since you have answer choices to chose from: 3|x − 3| + 2 = 14 3|-1-3| + 2 =14 3|-4| + 2 =14 3(4)+2 12 +2 = 14 14 = 14 3|x − 3| + 2 = 14 3|7 − 3| + 2 = 14 3|4| + 2 = 14 3(4) + 2 =14 12 + 2 =14 14 = 14 Since both -1 and 7 work, the solution to the equation is x= -1 or 7.

OpenStudy (coralskates):

thanks!!!!!!!!!!!!!!

OpenStudy (sushi121212):

You're welcome!